1. A business' return of investment will be allocated on three business partners; joey, james, and joe in the ratio 3:4:5, respectively. If the total income is P1,200,000, how much will james receive?


2. If dad wants to give P5,000 to his four children in the ratio 1:2:3:4 for their week allowance, how much will each of the four children receive?


3. If the tutorial fee is P600 per hour, how much should you pay for 5 hours?


4. Assuming they work at the same rate, how long will it take for 2 housekeepers to clean an entire house if it takes 4 days for 8 housekeepers to clean it?


5. Carla's birthday party will cost P3920 if she invites 14 guests. If the cost is directly proportional tto the number of invited guests, how much will it cost if she invites 56 guests?


1
Expert's answer
2021-11-12T01:20:14-0500

Total income = P1,200,000

  1. Ratio 3:4:5 Joey,James, Joe respectively

"Total ratio =3+4+5 =12"

James will receive "\frac{4}{12}×1,200,000 =P400,000"


2.

total amount is P5,000

Ratio = 1:2:3:4

total ratio = 1+2+3+4 = 10

Each of the four children will receive:

"\frac{1}{10}×5,000 = P500"


"\frac{2}{10}×5,000 =P1,000"


"\frac{3}{10}×5,000 = P1,500"


"\frac{4}{10}×5,000 =P2,000"


3.

If tutorial fee is P600 per hour,then for 5 hours it will be:

"\frac{5hr × P600}{1hr} =P3,000"


4.

8 housekeepers takes 4 days

2 housekeepers will take:

"\frac{8}{2}×4= 16days"


5.

birthday party will cost P3,920 is she invites 14 guests, if she invites 56 guests the cost will be:

"\frac{3920×56}{14}= P15,680"